OpenStudy (anonymous):

Can you check a little of my work, ill post a pic of them answers just make sure their right @Mertsj so I can give you a medal

OpenStudy (mertsj):

Post it and we'll see.

OpenStudy (mertsj):

The definition of cosine is side adjacent over hypotenuse. Look at your first problem. Which side is adjacent the angle theta?

OpenStudy (mathmale):

You might want to try entering 5/12 instead of the rounded off decimal fraction. First, plese look at the illustration and ask yourself what 5 and what 12 represent.

OpenStudy (anonymous):

adjacent is 5 and hypotenuse is 13

OpenStudy (mertsj):

Ok. The cavalry has arrived. Good luck.

OpenStudy (mathmale):

You two make a great team!

OpenStudy (anonymous):

Who?

OpenStudy (mathmale):

You and Mertsj. I mean that in the most positive sense.

OpenStudy (anonymous):

Lol oh okay, can you help me ?? I cant afford to get these wrong, which ones do I have right?

OpenStudy (mathmale):

the more you can show me your work, the more I'd be willing to give you feedback.

OpenStudy (anonymous):

Ok for the first one: 5 adjacent-12 opposite-13 hypot. .... so I did 5/13 because thats what you have to do for cosine

OpenStudy (mathmale):

True, but I'd much prefer that you remember and vocalize this as "the cosine of an angle is the ratio of the side of the triangle adjacent to the angle to the hypotenuse of the triangle," or just cos theta = adj/hyp.

OpenStudy (anonymous):

so the answer is .38?

OpenStudy (mathmale):

Have you any way of photographing your work and posting the image here? I'd prefer not to give you "right" or "wrong" answers, but yes, 5/13 does equal roughly 0.38.

OpenStudy (anonymous):

thats this one

OpenStudy (mathmale):

How many decimal places does your instructor want in answers?

OpenStudy (anonymous):

She doesnt say.

OpenStudy (mathmale):

What a nice illustration! Hey, Nic, please apply what you know about sin functions to find the sine of this angle theta.

OpenStudy (anonymous):

.92?

OpenStudy (mathmale):

Looks OK. Please type in the proper fraction that represents opp/hyp.

OpenStudy (anonymous):

12/13?

OpenStudy (mathmale):

Perfect. I'd prefer proper fraction answers because it shows me what you've done. Nice work. What is the tangent of this angle theta?

OpenStudy (anonymous):

12/5=2.4

OpenStudy (mathmale):

Great. Keep up the good work! Next question?

OpenStudy (anonymous):

Hold on let me solve it

OpenStudy (mathmale):

Not mandatory, but if you could type in your work, or take a photo of your work and share the image, I'd be happiest.

OpenStudy (mathmale):

I'm going to take a different tack here: I ask you to identify the values of the : opp side adj side hypotenuse formula for sine

OpenStudy (anonymous):

opp- 15 adj-8 hypot.-17 opp/hypot or 15/17=.88

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Nice work. Again, the more I see of your actual work, the more I can do for you in return. New problem?

OpenStudy (anonymous):

You're so helpful thank you so much and yes hold on

OpenStudy (anonymous):

I think this is right?

OpenStudy (mathmale):

Mind explaining your reasoning?

OpenStudy (mathmale):

Can you explain your reasoning using the word "similar"?

OpenStudy (anonymous):

Uh I am just going off another one I had like that, I dont really understand it?

OpenStudy (mathmale):

This is a strange math problem in that the original triangle is expanded/stretched, not "dilated' (UNLESS, of course, I have rocks in my head). If I give y ou a triangle and you make every side twice as long as before, the angles in question do not change, and the 2 triangles are called "similar." Have you heard of this before?

OpenStudy (anonymous):

oh okay yeah

OpenStudy (anonymous):

Its the same triangle so I am guessing it would still be 22/5 even if it was bigger?

OpenStudy (mathmale):

so my last comment pretty much answers this question. What's your conclusion? It's not the same triangle, but the 2 triangles are definitely "similar" in the mathematical sense, so yes, 22/5 is the tangent function value.

OpenStudy (anonymous):

Okay and I am lost on this one

OpenStudy (mathmale):

The tangent of an angle is the ratio of opp to adj, and while those may change as the triangle gets bigger or smaller, their ratio doesn't!

OpenStudy (mathmale):

Waiting for y our diagram to load; loading is very slow

OpenStudy (anonymous):

Okay no prob

OpenStudy (mathmale):

this question has to do with "cofunctions." The cosine is the "cofunction" of the sine. The principle involved here is that sin A = cos (90-A). So, sin 60 = cos (90-60) = cos 30.

OpenStudy (mathmale):

With this in mind, pleasae reread the qeustion anx see whther you can come up with the correct choice yourself. Forgive the misspellings; i'm aiming for speed more thn accuracy at the moment. :)

OpenStudy (anonymous):

30 ; 1?

OpenStudy (mathmale):

the proper name for this is "complementary angles."

OpenStudy (mathmale):

In the original problem, the given angle is 30 degrees. Have you subtracted that from 90 degrees? you need the "complement" of 30 degrees.

OpenStudy (anonymous):

60 degrees?

OpenStudy (mathmale):

right.

OpenStudy (anonymous):

and the second part is 1/2 ???

OpenStudy (mathmale):

sin 30 is 1/2, according to the problem statement. Therefore, cos ( ?? ) = 1/2

OpenStudy (mathmale):

See, the ratios (1/2 and 1/2) remain the same, but the function changes and the angle changes. sin 30 is 1/2, so cos ( ??? ) = 1/2.

OpenStudy (anonymous):

60 ; 1/2?

OpenStudy (mathmale):

You've already typed in the correct responses, but what the heck, please identify the missing angle again. sin 30 is 1/2, so cos ( ) is also 1/2.

OpenStudy (mathmale):

Right. perfect. Make sense now?

OpenStudy (anonymous):

Yessss rhank you so much!! I am going to takae the next assesment so I might need help or need u to check them

OpenStudy (mathmale):

Two angles that add up to 90 degrees are "complementary." Two that add up to 180 degrees are "supplementary." Youll see these also.

OpenStudy (mathmale):

Happy to do it! Just remember that i'm more interested in how you obtain your answers than I am in the answers themselves.

OpenStudy (anonymous):

Need help with this! Last one sorry @mathmale

OpenStudy (mathmale):

Nothing to be sorry about! You're doing well and i feel appreciated!

OpenStudy (mertsj):

@mathmale How do we know this isn't a test you are doing for someone?

OpenStudy (anonymous):

It is not a test I can confirm that!! I am doing them by myself, I only need help with the ones I really cannot figure out

OpenStudy (mathmale):

I'll let Nic speak for himself.

OpenStudy (anonymous):

I am a girl lol :-)

OpenStudy (anonymous):

Thanks again :))) @mathmale I ended up getting it

OpenStudy (mathmale):

I greatly enjoyed working with you! Keep up the good work! MM

OpenStudy (anonymous):

Thank you, Ill try